HIGH-RESOLUTION FOURIER-TRANFORM SPECTRA OF THE NO_{2} A–X ELCTRONIC BANDS AROUND 1 μm

\underline{J. ORPHAL}, S. VOIGT, S. DREHER, and J. P. BURROWS, Institute of Environmental Physics, University of Bremen, P. O. Box 330440, 28334 Bremen, Germany; R. JOST and A. DELON, Grenoble High Magnetic Field Laboratory, CNRS/MPI, BP 166 X, 38042 Grenoble Cedex 9, France.

The strong vibronic interactions between the A and X electronic states lead to a complex structure of the NO_{2} spectrum in the visible and near-IR. The long-wavelength tail of the spectrum is very weak due to small Franck-Condon factors. In 1965, Douglas and Huber^{a} recorded the first low-resolution spectra of the NO_{2} bands in the near infrared. In 1975, Brand et al.^{b} showed that ”hot” bands are important in this region. In the past, many high-resolution spectra of NO_{2} were recorded, however no rotationally resolved spectra of this important region were reported up to now. We present high-resolution spectra of the first vibronic bands of NO_{2} starting around 9740 cm^{−1}, and discuss vibrational assignments. The dense rotational structure is very difficult to assign, although accurate lower-state energies are available from high-resolution infrared spectra^{c}.

aA. E. Douglas and K. P. Huber, Can. J. Phys. 43, 74 (1965)

bJ. C. D. Brand, W. H. Chan, and J. L. Hardwick, J. Mol. Spectrosc. 56, 309 (1975)

cA. Perrin, C. Camy-Peyret, and J. M. Flaud, J. Quant. Spectrosc. Rad. Transfer 48, 645 (1992)